Please help me solve problem:
One number is twice the other. The sum of three times the smaller and two times the larger is 49. Find the numbers.

Answer by rapaljer(4551)

Let x = first number (the smaller)
2x = second number (the larger)

The equation is: "Three times the smaller plus two times the larger is 49."
In math symbols this is:
3( smaller) + 2(larger) = 49
3(x) + 2(2x) = 49
3x + 4x = 49
7x = 49
x = 7 smaller number
2x=14 larger number

Check:
Three times smaller plus two times larger is 49.
3(7) + 2(14) = 49
21 + 28 = 49 Check!

R^2 at SCC